# Fast nonparametric inference of network backbones for graph sparsification

### v1

@@ III.3 Comparison on Real Networks
Figure 7: Comparison of backbone metrics on real networks. The five backboning methods are compared using the same metrics as in Sec. III.2 on the real networks in the corpus described in Sec. III.3. In each of the first five panels we plot a matrix $\bm{B}(m)$ for a metric $m$ of interest such that $B_{ij}(m)$ is the fraction of real networks in the corpus for which $m_{i}\geq m_{j}$. High values $B_{ij}(m)$ indicate that the metric $m$ was higher for backboning method $i$ than backboning method $j$ in a large portion of real network instances. We plot the matrix $\bm{B}(m)$ for (a) the fraction of edges retained in the backbone, $m=E^{(b)}/E$; (b) the fraction of nodes with degree zero in the backbone; (c) the Hellinger distance (Eq. 17) between the strength distributions of the backbone and original network; (d) the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and (e) the relative reachability (Eq. 16) between the backbone and original graph. In panel (f) we plot the average network Jaccard similarity (Eq. 15) between the backbones generated by each pair of methods.
@@
We now compare the five backboning algorithms from Sec. III.2 on a large corpus of real network datasets coming from different application domains. We initially collected 69 networks from the Netzschleuder repository netz by retrieving all weighted networks with edge weights $w_{ij}\geq 1$ and less than $10^{7}$ edges. Non-integer-valued weights $w_{ij}$ were rounded to the nearest integer for analyses. As the methods in Sec. III.2 are all adaptable to both directed and undirected networks, we analyzed the 32 directed networks in the corpus using both edge directions separately (in-edges and out-edges) and analyzed the remaining 37 undirected networks with both edge directions simultaneously as described in Sec. II.1. Three directed networks and three undirected networks returned empty backbones for all five methods so were discarded from the results, leaving 58 directed and 34 undirected backbones constructed by each of the five methods. The networks in the final corpus represent a diverse array of domains, and (using the classification provided by netz ) in total we inferred 54 social network backbones, 11 informational network backbones, 9 biological network backbones, 7 technological network backbones, 6 transportation network backbones, and 5 economic network backbones using each of the five methods. The networks ranged in size from $E=91$ to $E=5,743,258$ with total weight ranging from $W=282$ to $W=2,541,576,441$ and average degree ranging from $\langle k\rangle=1.1$ to $\langle k\rangle=238.3$.
@@
Our first experiment with this real-world network corpus examines the extent to which the backbones differ between the global and local MDL backboning procedures described in Sec.s II.1 and II.2. In Fig. 6(a) we plot the fraction of the original edges retained in the local MDL backbone versus the fraction retained in the global MDL backbone for each of the backbones examined. We only include examples for which both methods found a non-empty backbone, noting the number of violating examples in the axis labels. We can see that the global MDL method tends to find empty backbones in a large fraction of the examples studied (25/92), indicating that this method is frequently not able to compress real network structure any better than a naive edge transmission. On the other hand, only two networks failed to be compressed by the more flexible local MDL method. When both methods give nontrivial backbones, we can see that they return a similar number of edges, with the local MDL method returning more edges in many cases but neither method exceeding $E^{(b)}/E\approx 0.35$. In Fig. 6(b) we plot the inverse compression ratios (Eq. 6) for each method, which indicates that when the global MDL method does compress it does so quite well, even compressing roughly $\approx 90\%$ of the information in the network relative to a naive edge transmission in some cases. We see that the global method compresses better than the local method in many cases, demonstrating that a global threshold can in fact provide an effective backbone for many real networks. Neither panel indicates any clear differentiation among the backbone structure of networks based on domain.
@@
In the next experiment we compare all five methods with respect to the metrics evaluated in Sec. III.2—specifically, the fraction of edges retained in the backbone; the fraction of nodes with degree zero in the backbone; the Hellinger distance (Eq. 17) between the strength distributions of the backbone and original network; the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and the relative reachability (Eq. 16) between the backbone and original graph. We omit analysis of the inverse compression ratio since trivially the global and local MDL methods will always obtain better compression than their global and local counterparts, and the direct comparison of the two MDL methods is done in Fig. 6. Fig. 7(a)-(e) shows the results of these experiments through pairwise comparisons among all pairs of backboning methods for each of the five measures. We can see that in the real networks the methods are roughly ordered like MDL $\text{MDL Local}\approx\text{Disparity}(E_{Local})>\text{MDL Global}>\text{Disparity}(\alpha=0.05)>\text{High Salience}$ when considering the number of edges, the total weight, the fraction of isolated nodes, and the relative reachability of the inferred backbones. The reverse pattern is observed for the Hellinger distance, with MDL Local and $\text{Disparity}(E_{Local})$ performing better than the other methods on a substantial fraction of networks. MDL Local has a slight edge over $\text{Disparity}(E_{Local})$ with respect to the fraction of isolated nodes and the reachability, while $\text{Disparity}(E_{Local})$ has a slight edge for the Hellinger distance.
@@
In Fig. 7(f) we plot the Jaccard similarity (Eq. 15) between the backbones generated by each pair of methods, averaged over all networks in the corpus. We observe very low levels of similarity among all pairs of backbones except for those generated by the MDL Local and $\text{Disparity}(E_{Local})$ methods, which have a moderately high average overlap of roughly $0.5$.
@@
Overall, the results in Sec. III.2 and Sec. III.3 indicate that the local MDL method and the Disparity Filter set to the same number of backbone edges as the local MDL method tend to best preserve the original network structure in many cases. Meanwhile, the global MDL method often finds empty backbones and does not compress relative to a naive transmission scheme, but in the instances that it does compress it performs well across the metrics while retaining fewer edges and providing even better compression than the local method. Setting the Disparity filter to the typical significance level of $\alpha=0.05$ tends to be too conservative in some real networks, suggesting that this parameter needs to be carefully tuned in practice for different applications. Finally, the High Salience Skeleton is quite conservative in its backbone estimation, as it highlights only a few links that are the most significant for global network routing.
@@ IV Conclusion
In this paper we develop a completely nonparametric framework for inferring the backbone of a weighted network which utilizes the minimum description length (MDL) principle to promote sparsity. Our method is adapted to infer both global and local network backbones, and is generalizable to any mixture of weight distributions over the backbone and non-backbone edges using a corresponding Bayesian generative model. We develop fast optimization schemes for our global and local MDL backboning objectives that are log-linear in the number of edges, allowing these principled methods to easily scale to networks with millions of edges. We compare our method with existing methods in a range of tasks on synthetic and real network data, showing that the proposed MDL methods are capable of substantially sparsifying a wide variety of networks while retaining meaningful global and local structural characteristics.
There are a number of potential avenues for future work extending our methods. As mentioned in Sec. II.3, our method is generalizable for inferring global and local network backbones under any mixture of weight distributions in the backbone and non-backbone. Here we study a model equivalent to a mixture of geometric distributions, but in future work it is important to examine other distributions which may provide better compression of networks with different levels of weight heterogeneity. One could also fully utilize the posterior distribution over backbones in this Bayesian framework to sparsify networks in a stochastic manner, similarly to some existing sparsification techniques spielman2008graph . Both of these extensions may incur a substantial additional computational burden for optimization and sampling compared to the fast greedy optimization performed here. One could also extend our framework to sparsify hypergraphs or other higher order networks with weight metadata. Finally, it may be possible to extend our method to infer sparse functional network backbones by exploiting regularities in dynamical information such as node or edge states.
@@ Acknowledgments
Appendix A Additional Comparisons on Synthetic Networks

### v3

@@ III.3 Comparison on Real Networks
Figure 7: Comparison of backbone metrics on real networks. The six backboning methods were compared using the same metrics as in Sec. III.2 on the real networks in the corpus described in Sec. III.3. In each of the first five panels we plot a matrix $\bm{B}(m)$ for a metric $m$ of interest such that $B_{ij}(m)$ is the fraction of real networks in the corpus for which $m_{i}\geq m_{j}$. High values $B_{ij}(m)$ indicate that the metric $m$ was higher for backboning method $i$ than backboning method $j$ in a large portion of real network instances. We plot the matrix $\bm{B}(m)$ for (a) the fraction of edges retained in the backbone, $m=E^{(b)}/E$; (b) the fraction of nodes with degree zero in the backbone; (c) the Hellinger distance (Eq. 23) between the strength distributions of the backbone and original network; (d) the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and (e) the relative reachability (Eq. 22) between the backbone and original graph. In panel (f) we plot the average network Jaccard similarity (Eq. 21) between the backbones generated by each pair of methods, with a different color scale to indicate the different interpretation of the heatmap.
@@
We now compare the six backboning algorithms from Sec. III.2 on a large corpus of real network datasets coming from different application domains. We initially collected 69 networks from the Netzschleuder repository netz by retrieving all weighted networks with edge weights $w_{ij}\geq 1$ and less than $10^{7}$ edges. Non-integer-valued weights $w_{ij}$ were rounded to the nearest integer for analyses since the microcanonical models were used. As the methods in Sec. III.2 are all adaptable to both directed and undirected networks, we analyzed the 32 directed networks in the corpus using both edge directions separately (in-edges and out-edges) and analyzed the remaining 37 undirected networks with both edge directions simultaneously as described in Sec. II.1. The networks in the final corpus represent a diverse array of domains, and (using the classification provided by netz ) in total we inferred 54 social network backbones, 11 informational network backbones, 9 biological network backbones, 7 technological network backbones, 6 transportation network backbones, and 5 economic network backbones using each of the six methods. The networks ranged in size from $E=91$ to $E=5,743,258$ with total weight ranging from $W=282$ to $W=2,541,576,441$ and average degree ranging from $\langle k\rangle=1.1$ to $\langle k\rangle=238.3$.
@@
Our first experiment with this real-world network corpus examines the extent to which the backbones differ between the global and local MDL backboning procedures described in Sec.s II.2. In Fig. 6(a) we plot the fraction of the original edges retained in the local MDL backbone versus the fraction retained in the global MDL backbone for each of the backbones examined. We only include examples for which both methods found a non-empty backbone, noting the number of violating examples in the axis labels. We can see that the global MDL method tends to find empty backbones in a large fraction of the examples studied (25/92), indicating that this method is frequently not able to compress real network structure any better than a naïve edge transmission. On the other hand, only two networks failed to be compressed by the more flexible local MDL method. When both methods give nontrivial backbones, we can see that they return a similar number of edges, with the local MDL method returning more edges in many cases but neither method exceeding $E^{(b)}/E\approx 0.35$. In Fig. 6(b) we plot the inverse compression ratios (Eq. 20) for each method, which indicates that when the global MDL method does compress it does so quite well, even compressing roughly $\approx 90\%$ of the information in the network relative to a naive edge transmission in some cases. We see that the global method compresses better than the local method in many cases, demonstrating that a global threshold can in fact provide an effective backbone for many real networks. Neither panel indicates any clear differentiation among the backbone structure of networks based on domain.
@@
In the next experiment we compare all six backboning methods of Sec. III.2 with respect to the same metrics—specifically, the fraction of edges retained in the backbone; the fraction of nodes with degree zero in the backbone; the Hellinger distance (Eq. 23) between the strength distributions of the backbone and original network; the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and the relative reachability (Eq. 22) between the backbone and original graph. We omit analysis of the inverse compression ratio since trivially the global and local MDL methods will always obtain better compression than their global and local counterparts, and the direct comparison of the two MDL methods is done in Fig. 6. Fig. 7(a)-(e) shows the results of these experiments through pairwise comparisons among all pairs of backboning methods for each of the six measures. We can see that in the real networks the methods are roughly ordered like Percolation>MDL $\text{Percolation}>\text{MDL Local}\approx\text{Disparity}(E_{Local})>\text{MDL Global}>\text{Disparity}(\alpha=0.05)>\text{High Salience}$ when considering the number of edges, the total weight, the fraction of isolated nodes, and the relative reachability of the inferred backbones. Meanwhile, the reverse pattern is observed for the Hellinger distance. MDL Local has a slight edge over $\text{Disparity}(E_{Local})$ with respect to the fraction of isolated nodes and the reachability, while $\text{Disparity}(E_{Local})$ has a slight edge for the Hellinger distance. The percolation backbone performs the best overall in these tests but at the cost of retaining a substantial fraction of edges (an average of $84\%$!).
@@
In Fig. 7(f) we plot the Jaccard similarity (Eq. 21) between the backbones generated by each pair of methods, averaged over all networks in the corpus. We observe very low levels of similarity among all pairs of backbones except for those generated by the MDL Local and $\text{Disparity}(E_{\text{Local}})$ methods, which have a moderately high average overlap of roughly $0.5$.
@@
Overall, the results in Sec. III.2 and Sec. III.3 indicate that the local MDL method and the Disparity Filter set to the same number of backbone edges as the local MDL method tend to best preserve the original network structure in many cases. Meanwhile, the global MDL method often finds empty backbones and does not compress relative to a naive transmission scheme, but in the instances that it does compress it performs well across the metrics while retaining fewer edges and providing even better compression than the local method. Setting the Disparity filter to the typical significance level of $\alpha=0.05$ tends to be too conservative in some real networks, suggesting that this parameter needs to be carefully tuned in practice for different applications. The High Salience Skeleton is quite conservative in its backbone estimation, as it highlights only a few links that are the most significant for global network routing. And finally, the percolation backbone tends to preserve connectivity and weight heterogeneity but does not provide substantial sparsification, retaining a majority of the edges in most networks.
III.4 Comparison of Spreading Dynamics on Network Backbones
Figure 8: Speeding up epidemic calculations with network backbones. Fractional size of the epidemic cluster (Eq. 26) versus the per-contact transmission probability $p$ for the message passing system of Eq.s 24 and 25. The percolation threshold estimated using Eq. 27 for each network backbone and the full network (black) are shown with dotted vertical lines. The average error in $S$, runtime savings, and error in the percolation threshold (as described in Sec. III.4) are shown in the right panel.
As an example downstream application of our backboning methods, we examine their performance against baselines for sparsifying networks while retaining dynamical properties relevant for spreading and percolation calculations. Specifically, we adapt a standard method for estimating the epidemic threshold from long-term transmission probabilities among each pair of nodes newman2018networks to the weighted network case, and study how accurate and computationally expensive each backbone is for this calculation on an example network.
In the original model discussed in newman2018networks , there is a single parameter $\phi$ characterizing the probability that a node $i$ has strong enough contact with any other node $j$ for the disease to get passed from $j\to i$ if $j$ ends up with the disease. This is analogous to the occupation probability in a bond percolation process. In our weighted network adaptation, we consider the parameter $\phi(w_{ij})$ characterizing this probability to be a monotonic function of the weight $w_{ij}$ of the edge between nodes $i$ and $j$. (For simplicity, we restrict our analyses to undirected networks.) This mechanism corresponds to $i$ and $j$ having a greater connectivity and potential to transmit the disease to each other when they have a higher weight edge between them. Specifically, we consider a process in which the weight $w_{ij}$ can be interpreted as the frequency of contact between $i$ and $j$, and that there is a probability $p$ for the disease to spread independently from each contact, so that $\phi(w_{ij})=1-(1-p)^{w_{ij}}$.
By varying $p\in[0,1]$, we can find a percolation transition in this model in which an extensive giant component forms and an epidemic can spread to the entire network. We can compute this threshold using a method analogous to that of karrer2014percolation in which a constant occupation probability is considered. Letting $u_{i\to j}$ be the probability that node $i$ is not connected to giant component through node $j$—in other words, the probability that $j$ does not contribute to the epidemic through passing the disease to node $i$—we have
$u_{i\to j}=1-\phi(w_{ij})+\phi(w_{ij})\prod_{k\in\mathcal{N}_{j\setminus i}}u_{j\to k},$ (24)
where $\mathcal{N}_{j\setminus i}$ is the set of nodes attached to $j$ other than node $i$. The probability that $i$ is part of the giant cluster (e.g. epidemic cluster) is then given by
$S_{i}=1-\prod_{j\in\mathcal{N}_{i}}u_{i\to j},$ (25)
where $\mathcal{N}_{i}$ is the set of nodes connected to $i$, and the fractional size $S$ of the giant cluster is
$S=\frac{1}{N}\sum_{i=1}^{N}s_{i}=1-\frac{1}{N}\sum_{i=1}^{N}\prod_{j\in\mathcal{N}_{i}}u_{i\to j}.$ (26)
Technically, the above calculation is only exact on trees—otherwise, we cannot assume the independence of $j$’s neighbors to write $\prod_{k\in\mathcal{N}_{j\setminus i}}u_{j\to k}$. However, it often gives a good approximation in practice for tree-like sparse networks. (If one is interested in more complex approximations accounting for dependence among node neighbors yedidia2003understanding ; kirkley2021belief , sparsifying the networks ahead of time through backboning may be necessary, for tractability.)
Equations 24 and 25 define a system of message passing equations that can be used to solve for the size of the percolating epidemic cluster of any network $G$ as a function of the per-contact transmission probability $p$. To do this, the probabilities $u_{i\to j}$ are initialized to random values, and Eq. 24 is iterated until convergence, at which point $S_{i}$ can be computed for each node $i$ using Eq. 25.
Fig. 8 shows the results of simulating this message passing system on the reality mining network of student contacts collected from mobile phone data in eagle2006reality , for $p$ ranging from $10^{-6}$ to $1$. The system was simulated on the full network (black curve) as well as the backbones constructed using the six methods of Sec. III.2. To compare these curves, we compute the average absolute error $\langle|S-S_{0}|\rangle$ between the values $S$ associated with each backboning method and the full network’s values $S_{0}$. We can observe that the percolation-based backboning method performs the best with respect to this metric, although we will see that it does not provide a substantial computational speedup since it retains a large fraction of the edges. The local MDL method and the disparity filter with $E=E_{\text{Local}}$ perform simililarly under this measure, still providing a good approximation of the full network’s giant component size at each $p$, while the global MDL method performs slightly worse, deviating from the desired curve for high values of $p$. The disparity filter with $\alpha=0.05$ performs worse yet, with noticeably large deviations as the giant component grows in size, and the high salience skeleton—with very little retained in the backbone—performs the worst out of the six methods for this task.
We can notice that at a critical probability $p^{(c)}\in[0,1]$, the system transitions from a fragmented state with many small connected components ($S\approx 0$) to having a single extensive giant connected component ($S\approx 1$). There is a trivial fixed point in the message passing system (Eq. 24) of $u_{i\to j}=1$—in other words, when no one is in percolating cluster ($S=0$). We can therefore find the percolation transition by expanding in $u_{i\to j}=1-\epsilon_{i\to j}$ for small $\epsilon_{i\to j}$ and identifying at which value of $p$ this fixed point is no longer stable. When the trivial fixed point is stable, we have no giant component, and when it is unstable, we have a giant component. (A similar calculation is performed in karrer2014percolation for constant occupation probability $\phi$.)
Plugging in $u_{i\to j}=1-\epsilon_{i\to j}$ to Eq. 24 and expanding for small $\epsilon_{i\to j}$ gives
$\epsilon_{i\to j}=\phi(w_{ij})\sum_{k\in\partial_{j\setminus i}}\epsilon_{j\to k}=\sum_{i\to j,l\to k}B^{(\phi)}_{i\to j,l\to k}\epsilon_{l\to k},$ (27)
where $B^{(\phi)}_{i\to j,l\to k}=\phi(w_{ij})\delta_{jl}(1-\delta_{ik})$ are entries—indexed by edges $i\to j$ and $l\to k$—in a modified version of the non-backtracking (or Hashimoto) matrix that account for the transmission probabilities. The system will percolate when the above fixed point is unstable, i.e. when $\bm{B}^{(\phi)}$ has a leading eigenvalue of magnitude greater than $1$. We can then identify the value $p^{(c)}$ numerically without the need for extensive message passing simulations by evaluating the leading eigenvalue of $\bm{B}^{(\phi)}$ at some starting guess $p$ and running a binary search over $p$ until the leading eigenvalue of $\bm{B}^{(\phi)}$ has magnitude close to $1$. The value at which this terminates will be an estimate of $p^{(c)}$.
In Fig. 8 we can examine how well each backboning method preserves the percolation threshold $p^{(c)}$ of the full network and how much runtime it saves. A good balance of these two factors is ideal in practice, as we desire accuracy while saving computational cost. (For the particular example network studied here, we do not face any computational bottleneck. But this serves as a proof-of-concept for application to much larger networks.) We measure the runtime savings $\langle t/t_{0}\rangle$ as the fractional reduction relative to the runtime $t_{0}$ for the full network, averaged over all evaluations of the binary search for $p^{(c)}$. We then measure the error in the percolation threshold as $|p^{(c)}-p^{(c)}_{0}|$, where $p^{(c)}_{0}$ is the threshold for the full network.
We can see that, while the percolation-based backbone has the most accurate estimation of $p^{(c)}$, it fails to reduce the runtime of the computation substantially, requiring on average $74\%$ of the runtime of the full network calculation to compute the leading eigenvalues. On the opposite extreme, the leading eigenvalue calculations on the high salience skeleton require almost no runtime, but are completely inaccurate and estimate $p^{(c)}=1$. The disparity filter with $\alpha=0.05$ has the next highest error in $p^{(c)}$, though provides runtimes much lower than other methods (other than the high salience skeleton). The global MDL, local MDL, and disparity filter set to $E=E_{\text{Local}}$ find a good middle ground, allowing for a roughly 10×10$\times$10 × speedup in the calculation while approximating the percolation transition to within $6$ decimal places.
These results together indicate that the MDL backbones are capable of preserving not only structural, but also dynamical properties of networks, despite not explicitly involving dynamics in their formulation.
@@ IV Conclusion
In this paper we develop a completely nonparametric framework for inferring the backbone of a weighted network which utilizes the minimum description length (MDL) principle to promote sparsity. Our method is adapted to infer both global and local network backbones, and is generalizable to any mixture of weight distributions over the backbone and non-backbone edges using a flexible Bayesian generative model. We develop fast exact optimization schemes for our global and local MDL backboning objectives that are log-linear in the number of edges, allowing these principled methods to easily scale to networks with millions of edges. We compare our method with existing methods in a range of tasks on synthetic and real network data, showing that the proposed MDL methods are capable of substantially sparsifying a wide variety of networks while retaining meaningful global and local structural characteristics.
There are a number of potential avenues for future work extending our methods. One critical limitation of the proposed method is that it does not apply to unweighted graphs, since the weights are used for compression to identify a meaningful backbone. Extending the MDL framework for backboning unweighted graphs is therefore an important avenue for future work. Additionally, as mentioned in Sec. II.1, our method is generalizable for inferring global and local network backbones under any mixture of weight distributions in the backbone and non-backbone, with a subset of these permitting exact minimization with Algorithms 1 and 2. Here we study a microcanonical model asymptotically equivalent to a canonical model with geometric weights, but in future work it is important to examine other distributions which may provide better compression of networks with different types of weight heterogeneity. One could also fully utilize the posterior distribution over backbones in this Bayesian framework to sparsify networks in a stochastic manner, similarly to some existing sparsification techniques spielman2008graph . Both of these extensions may incur a substantial additional computational burden for optimization and sampling compared to the fast greedy optimization performed here, so present new challenges for practical implementation. One could also extend our framework to sparsify hypergraphs or other higher order networks with weight metadata. Finally, it may be possible to extend our method to infer sparse functional network backbones by exploiting regularities in dynamical information such as node or edge states.
@@ Acknowledgments
Appendix A Asymptotic Equivalence Between Microcanonical Model and Canonical Model with Geometrically Distributed Weights
In this Appendix we demonstrate the approximate equivalence between the microcanonical model of Sec. II.2 and the canonical model of Sec. II.1 when weights are distributed according to a geometric distribution. Letting $P(w|\theta)=\theta(1-\theta)^{w-1}$ be a geometric distribution and $P(\theta)=1$ a uniform prior on $[0,1]$ (which is conjugate to $P(w|\theta)$), we have that Eq. II.1 can be written as
$\mathcal{L}$ (${}_{C}^{(\text{global})}(G^{(b)})=\log(E+1)+\log{E\choose E^{(b)}}-\log\int\prod_{e\in G^{(b)}}\theta_{1}(1-\theta_{1})^{w_{e}-1}d\theta_{1}-\log\int\prod_{e\in G\setminus G^{(b)}}\theta_{0}(1-\theta_{0})^{w_{e}-1}d\theta_{0}$
[…]
$=\log(E+1)+\log{E\choose E^{(b)}}-\log B(E^{(b)}+1,W^{(b)}-E^{(b)}+1)-\log B(E-E^{(b)}+1,(W-W^{(b)})-(E-E^{(b)})+1)$
$=\log(E+1)+\log(W^{(b)}+1)+\log(W-W^{(b)}+1)+\log{E\choose E^{(b)}}+\log{W^{(b)}\choose E^{(b)}}+\log{W-W^{(b)}\choose E-E^{(b)}}$ (28)
$=\mathcal{L}_{M}^{(\text{global})}(G^{(b)})+\delta,$
where
$\delta=\log\frac{(W^{(b)}+1)(W-W^{(b)}+1)(W^{(b)})(W-W^{(b)})}{(W-E+1)(E^{(b)})(E-E^{(b)})}\leq\log\frac{(W/2+1)^{2}(W/2)^{2}}{(W-E+1)(E-1)}$ (29)
for $E^{(b)}\in[1,E/2]$. For $E^{(b)}=0$, we only evaluate the second integral since there are no edges in the backbone $G^{(b)}$, and the resulting expressions differ by $\delta=\frac{(W+1)W}{(W-E+1)E}$. Regardless, the correction scales like $\delta\sim O(\log W+\log E)$, and so we have
$\mathcal{L}_{C}^{(\text{global})}(G^{(b)})=\mathcal{L}_{M}^{(\text{global})}(G^{(b)})+O(\log W+\log E),$ (30)
and the canonical and microcanonical expressions are asymptotically equal for $W,E\gg 1$.
Similarly, the microcanonical neighborhood-level description length of Eq. II.1 is equivalent to the neighborhood-level description length of Eq. II.2 up to a correction of $O(\log s_{i}+\log k_{i})$, making these objectives asymptotically equivalent for $s_{i},k_{i}\gg 1$. However, the full canonical local description length of Eq. 10 is only equivalent to the corresponding microcanonical expression in Eq. 14 up to $-\log P(\bm{s})=\log{N+W-E-1\choose W-E}$ in addition to a correction of $O(N\langle\log k+\log s\rangle)$ which is of a similar size. Despite this discrepancy in the network-level description lengths for the local backboning methods, since the backboning is done for each neighborhood $\partial_{i}$ separately the microcanonical model is easily capable of inferring planted backbone structure generated from the canonical model (as seen in Fig. 2).
Appendix B Greedy Optimality in Numerical Experiments
Here we plot the results of reconstruction experiments using the same generative model as in Sec. III.1, except with small enough networks ($N=6$) to allow for exhaustive enumeration over all backbones $G^{(b)}\subseteq G$ to verify the optimality of Algorithm 1 and Algorithm 2. Figure 9 shows the compression ratio (Eq. 20), obtained using both the greedy method (x-axis) and exact enumeration over possible backbones (y-axis), for all experimental trials shown in Fig. 2. We find that the greedy algorithm identifies an identical description length to exact enumeration in all cases up to machine precision.
Figure 9: Optimality of greedy algorithm. Inverse compression ratios for all reconstruction experiments in Fig. 2, performed on networks with $N=6$ and $k\in\{2,3\}$. The results obtained from the greedy methods of Algorithm 1 and Algorithm 2 are shown on the x-axis, while the results obtained through exhaustive enumeration over possible backbones are shown on the y-axis. We find that the greedy algorithms give identical results to exact enumeration, as expected, for all cases studied.
Appendix C Additional Comparisons on Synthetic Networks